OpenStudy (anonymous):

\[(x-1)y''+y'=0\] \[y=\sum_{n=0}^{\infty} a_n x^n\] \[y'=\sum_{n=1}^{\infty} a_n x^{n-1}\] \[y''=\sum_{n=2}^{\infty} n(n-1)a_n x^{n-2}\]

OpenStudy (anonymous):

\[xy''-y''+y'=0\] \[x\sum_{n=2}^{\infty}n(n-1)a_nx^{n-2}-\sum_{n=2}^{\infty}n(n-1)a_nx^{n-2}+\sum_{n=1}^{\infty} a_n x^{n-1}=0\]